# SWIFT: Mapping Sub-series with Wavelet Decomposition Improves Time Series Forecasting

### v1

@@ 1 Introduction
Figure 1: Performance of Mean Squared Error (MSE) on a simple synthetic non-stationary signal. More details about the experimental settings can be found in Appendix.
@@ 2.1 Efficient linear forecasters
Since (Zeng et al., 2023) shows that a simple one-layer linear model can outperforms Transformer forecasters (Zhou et al., 2021; Wu et al., 2021; Zhou et al., 2022) in almost all cases, there has been a rapid emergence of linear forecasters (Oreshkin et al., 2020; Das et al., 2023; Liu et al., 2023) in LTSF. The impressive performance and efficiency continuously challenge this direction.
Recently, FITS (Xu et al., 2024) introduced a frequency-domain interpolation strategy that utilizes low-pass filters and FFT (Brigham and Morrow, 1967) for time series modeling. By maintaining a parameter scale at $10k$ level, FITS surpassed several existing Transformer models in both inference speed and forecasting performance, establishing itself as a benchmark in the field.
However, FFT assumes that signals are stationary, limiting its ability to capture the temporal localization of transient or non-stationary signals (Liu et al., 2022). Furthermore, its approach of applying a global frequency domain transformation to the entire signal often results in suboptimal performance when dealing with strong boundary effects. Meanwhile, the limited representational capacity of a single-layer linear model typically necessitates a longer look-back window to prevent underfitting and distribution shifts. The parameter count of FITS is primarily determined by the length of the look-back window due to its interpolation-based prediction approach. Consequently, The efficiency of FITS decreases significantly as the lookback window increases.
@@ 2.2 DWT method
As a powerful method for time-frequency analysis, DWT is widely used in tasks dealing with time series. (Yang et al., 2022) decomposed the time-series using wavelet decomposition and then utilized CNN and LSTM for prediction. FEDformer (Zhou et al., 2022) combines Wavelet Transform with frequency enhanced strategy and attention mechanism to capture long range dependencies. (Sasal et al., 2022) utilize a maximal overlap discrete wavelet transformation and build a local transformer model for time-series forecasting. Besides, DWT is often used for anomaly detection. For instance, (Bhattacharya et al., 2022) used wavelet transform for signal denoising and damage localization. Recently, (Arabi et al., 2024) proposed a method for data augmentation in time-series prediction tasks, which is used to obtain more diverse sequences by eliminating or swapping wavelet coefficients.
@@ 3.1 LTSF problem definition
In multivariate LTSF, time series data contain multiple variables or channels at each time step. Given historical values $\mathcal{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{L_{x}}~{}|~{}\mathbf{x}_{i}\in\mathbb{R}^{d}\}$ where $d$ represents the number of variables and $L_{x}$ represents the length of the look-back window, the objective of LTSF is to predict future values $\mathcal{Y}=\{\mathbf{y}_{1},\ldots,\mathbf{y}_{L_{y}}~{}|~{}\mathbf{y}_{i}\in\mathbb{R}^{d}\}$. The output length $L_{y}$ is usually much longer than the length of the lookback window $L_{x}$, and the feature dimension is not limited to the univariate case ($d\geq 1$).
@@ 3.2 DWT and time-frequency domain
We select Haar because it can make the transform fast and stable, which increases the speed of reasoning across our framework.
@@ 4.1 Structure Overview
We propose the SWIFT which is shown in Figure 2. Firstly, the time-series is decomposed by $1_{st}$ order DWT. Then the high-frequency component and the low-frequency component are concatenated and mapped using the same linear layer after passing through a learnable filter which is a convolutional layer. Finally, the prediction is obtained by performing IDWT on the new components obtained from the mapping.
Figure 2: Overall structure of SWIFT. (i)The DWT module decomposes the time series into two sub-series: the approximation coefficient and the detail coefficient, based on the Haar wavelet; (ii) The convolutional layer is applied for filtering and feature aggregation. (iii)Linear or MLP is used for the mapping of sub-series to make prediction.
@@ 4.2 SWIFT Components
After obtaining the representation $\mathbf{Y}$, SWIFT extracts information from this representation by means of convolution and Mapping. We will elaborate on this in section 4.2 and section 4.2.
@@ 5.1 Forecasting results
We carefully choose well-acknowledged forecasting models as our benchmark, including (1) Transformer-based methods: FEDformer (Zhou et al., 2022), PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2024). (2) Efficient Linear-based methods: DLinear (Zeng et al., 2023), FITS (Xu et al., 2024), CycleNet (Lin et al., 2024). (3): TCN-based methods: TimesNet (Wu et al., 2022). we use prediction lengths of 96, 192, 336, 720 and set the histor-ical horizon length to T= 720 for our model. For other baseline models, we follow (Xu et al., 2024) by treating the historical horizon T as a hyperparameter and using the settings from their original papers, since models may exhibit performance degradation with increasing historical horizons. For example, the performance of iTransformer on the ETT datasets declines as the length increases. As for CycleNet, We use a lookback window length of 720 because (Lin et al., 2024) shows that this length yields the best prediction result. We rerun all the experiment with code and script provided by their official implementation.
@@
Our method is trained with the ADAM optimizer (Kingma and Ba, 2015), using OneCycleLR strategy (Paszke et al., 2019) to adjust the learning rate. We fix the length of lookback window as 720, and we set the predicted window among {96, 192, 336, 720}. As for the kernel size of the convolution layer, we choose suitable size in {3, 9, 13, 17}.
@@
Comprehensive forecasting results are listed in Table 1 and Table 2 with the best in bold and the second underlined. The lower MSE indicates the more accurate prediction result. As shown in table 1 and table 2, SWIFT performs well in the forecasting task. Compared to FITS, the most powerful and efficient model currently, SWIFT achieves comparable or even superior performance in all 7 datasets.
Table 3 presents the number of trainable parameters and MACs for various linear-based time series forecasting (TSF) models using a look-back window of 720 and a forecasting horizon of 96 on the Electricity dataset. The table clearly demonstrates the exceptional efficiency of SWIFT compared to other models. Among all efficient models, SWIFT stands out with significantly fewer parameters and much faster training times. SWIFT-Linear requires only 15% of the FITS parameters and 60% of its MACs, while achieving comparable or even superior performance to these state-of-the-art efficient forecasting models. It should be noted that SWIFT-Linear’s parameter count is also much lower than Dlinear, which has 139.7K parameters. Moreover, while FITS’ parameter count increases rapidly when using long look-back windows for forecasting, SWIFT does not exhibit this issue.
@@
Table 3: Number of trainable parameters, MACs, and training time of models with less than 1M parameters, under look-back window=720 and forecasting horizon=96 on the Electricity dataset.
Ablations
@@
Table 4: Ablation study 1 of SWIFT. For dataset etth1 and ettm1, we choose Linear version of SWIFT; for dataset traffic, we use MLP version of SWIFT
Table 5: Ablation study 2 of SWIFT. IMP is the improvement between with $Share$ and $Split$ result, where a larger value indicates a better improvement. $Share$ corresponds to the use of only one linear layer, while $Split$ corresponds to the use of one linear layer for each frequency band. The model we used is SWIFT / Linear
@@
$W_{s}=\beta_{l}W_{l}+\beta_{h}W_{h}+\epsilon$
@@ 6 Conclusion and Future work
Appendix A Experimental Details
A.1 Datasets
We evaluate the performance of our proposed SWIFT on 7 popular datasets, including Weather, Traffic, Electricity, and ETT datasets. The Weather dataset, including 21 meteorological indicators such as air temperature and humidity, is collected every 10 minutes from the Weather Station of the Max Planck Institute for Biogeochemistry in 2020. The Traffic Wu et al. [2021] dataset contains hourly traffic data measured by 862 sensors on San Francisco Bay area freeways, which has been collected since January 1, 2015. The Electricity dataset collects the hourly electricity consumption of 321 clients from 2012 to 2014. The ETT (Electricity Transformer Temperature) datasets contain two visions of the sub-dataset: ETTh and ETTm, collected from electricity transformers every 15 minutes and 1 hour between July 2016 and July 2018. Thus, in total we have 4 ETT datasets (ETTm1, ETTm2, ETTh1, and ETTh2) recording 7 features such as load and oil temperature. The details about these datasets are summarized in Table 7.
Datasets ETTh1 ETTh2 ETTm1 ETTm2 Electricity Weather Traffic Exchange
Variables 7 7 7 7 321 21 862 8
Timesteps 17420 17420 69680 69680 26304 52696 17544 7588
Frequency Hourly Hourly 15min 15min Hourly 10min Hourly Daily
Information Electricity Electricity Electricity Electricity Electricity Weather Traffic Economy
Table 7: The details of datasets.
A.2 Baselines
We choose several well-acknowledged and state-of-the-art models for comparison to evaluate the effectiveness of our proposed SWIFT for time series forecasting, including MLP-based models and Transformer-based models. We introduce these models as follows:
FITS Xu et al. [2024] performs time series analysis through interpolation in the complex frequency domain, enjoying low cost with 10k parameters.
TimesNet Wu et al. [2022] transforms 1D time series into a set of 2D tensors based on multiple periods to analyse temporal variations. The above transformation allows the 2D-variations to be easily captured by 2D kernels with encoding the intraperiod- and interperiod-variations into the columns and rows of the 2D tensors respectively.
DLinear Zeng et al. [2023] utilizes a simple yet effective one-layer linear model to capture temporal relationships between input and output sequences.
Informer Zhou et al. [2021] enhances Transformer with KL-divergence based ProbSparse attention for $O(L\operatorname{log}L)$ complexity, efficiently encoding dependencies among variables and introducing a novel architecture with a DMS forecasting strategy.
FEDformer Zhou et al. [2022] implements sparse attention with low-rank approximation in frequency domain, enjoying linear computational complexity and memory cost. And it proposes mixture of experts decomposition to control the distribution shifting.
PatchTST Nie et al. [2023] divides time series data into subseries-level patches to extract local semantic information and adopts channel-independence strategy where each channel shares the same embedding and Transformer weights across all the series.
iTransformer Liu et al. [2024] inverts the structure of Transformer without modifying any existing modules by encoding individual series into variate tokens. These tokens are utilized by the attention mechanism to capture multivariate correlations and FFNs are adopted for each variate token to learn nonlinear representations.
Appendix B Visualizations
We applied Discrete Wavelet Transform (DWT) using the Haar wavelet to three datasets, decomposing each into low-frequency (approximation) and high-frequency (detail) components, as shown in Figure 7, Figure 7 and Figure 7. The low-frequency component reduces data length by half, enhancing processing efficiency while preserving key patterns like trends and structural features. The high-frequency component captures local fluctuations and fine-grained variations, crucial for detailed data analysis.
DWT enables efficient data compression and simplification while retaining core information. The low-frequency component optimizes the SWIFT model’s inference efficiency, reducing parameters and computational complexity, while the high-frequency component ensures detailed information integrity, supporting predictive performance. This approach provides a robust foundation for time series analysis and prediction, balancing computational efficiency and information preservation for real-time applications.
Figure 5: Visualization of DWT decomposition on the Traffic.
Figure 6: Visualization of DWT decomposition on the Weather.
Figure 7: Visualization of DWT decomposition on the ETT.
Models SWIFT FITS TimesNet Anomaly THOC Omni Stationary DGHL OCSVM IForest LightTS Dlinear
SMD 99.92 99.95 85.81 92.33 84.99 85.22 84.72 N/A 56.19 53.64 82.53 77.1
PSM 96.47 93.96 97.47 97.89 98.54 80.83 97.29 N/A 70.67 83.48 97.15 93.55
SMAP 68.16 70.74 71.52 96.69 90.68 86.92 71.09 96.38 56.34 55.53 69.21 69.26
MSL 58.51 78.12 85.15 93.59 89.69 87.67 77.5 94.08 70.82 66.45 78.95 84.88
Table 8: Anomaly detection result of F1-scores on 4 datasets. The best result is highlighted in bold, and the second best is highlighted with underline.
Appendix C Anomaly detection results
Datasets
We conducted experiments on 4 benchmark datasets used by [Xu et al., 2022]: SMD (Server Machine Dataset), PSM (Polled Server Metrics), MSL (Mars Science Laboratory rover) and SMAP (Soil Moisture Active Passive satellite).
Baseline
We compare SWIFT with models such as TimesNet [Wu et al., 2022], Anomaly Transformer [Xu et al., 2022], THOC [Shen et al., 2020], Omnianomaly [Su et al., 2019], DGHL [Challu et al., 2022]. Following TimesNet, we also compare the anomaly detection performance with other models [Zeng et al., 2023].
Main results
Table 8 demonstrates the performance of SWIFT on various datasets. SWIFT achieves near-perfect F1 scores of 99.92% and 96.487% on the SMD and PSM datasets, respectively, and achieves about 2.5% outperformance on PSM compared to FITS, demonstrating its high-precision performance in anomaly detection. By introducing the time-frequency domain, SWIFT can easily identify the anomalies that are difficult to recognize in the time domain, including the anomalies that introduce unexpected time-frequency components. In contrast, models such as TimesNet, Anomaly Transformer and Stationary Transformer do not perform as well as SWIFT on these datasets.
However, SWIFT performs relatively low on the SMAP and MSL datasets. Due to the binary event data nature of these datasets, time-frequency feature representations may have difficulty capturing these features effectively. In this case, time-domain modeling is more desirable because the raw data format is already compact enough. As a result, models designed for anomaly detection, such as THOC and Omni Anomaly, achieve higher F1 scores on these datasets.
It is worth noting that SWIFT has an extremely low number of parameters ranging from 0.625k-2.5k for the anomaly detection task. This feature allows SWIFT to be deployed on almost any edge device. In addition, SWIFT’s inference speed is impressive, reaching sub-millisecond levels, much faster than the latency associated with larger models or communication overheads. This speed underscores SWIFT’s suitability as a first-response tool for rapid detection of critical errors.
Appendix D Implementation details
Non-staionay data
We built a simulated non-stationary dataset and found that existing models cannot perform well in predicting non-stationary series. We give the code to build the Dataset here.
{minted}
[fontsize=, linenos, frame=lines, breaklines]python class NonStationaryTimeSeriesDataset(Dataset): def __init__(self, seq_length_input, seq_length_output, frequencies, amp_range, total_length, is_test=False): ””” Args: seq_length_input: Length of input sequences. seq_length_output: Length of output sequences. frequencies: List of frequency components for signal generation. amp_range: Amplitude range for signal generation. total_length: Total length of the time series. is_test: Whether the dataset is for testing (to introduce distinct temporal patterns). ””” self.seq_length_input = seq_length_input self.seq_length_output = seq_length_output self.frequencies = frequencies self.amp_range = amp_range self.total_length = total_length self.is_test = is_test
# Generate one long continuous sequence self.sequence = self._generate_sequence()
def _generate_sequence(self): # Time axis t = np.linspace(0, 2, self.total_length) # Normalize time axis to [0, 2]
# Generate signal with multiple frequency components amplitude = np.random.uniform(*self.amp_range) phase = np.random.uniform(0, 2 * np.pi, len(self.frequencies)) # Different phase for each frequency base_signal = sum( amplitude * np.sin(2 * np.pi * freq * t + phi) for freq, phi in zip(self.frequencies, phase) )
# Generate non-stationary signal: time-dependent frequency modulation if self.is_test: non_stationary_signal = sum( amplitude * np.sin(2 * np.pi * freq * t**2 + phi) for freq, phi in zip(self.frequencies, phase) ) else: non_stationary_signal = sum( amplitude * np.sin(2 * np.pi * freq * (t + 0.2 * np.sin(0.5 * t)) + phi) for freq, phi in zip(self.frequencies, phase) )
# Combine signals with trends and random noise trend = np.sin(0.5 * t) + 0.2 * t # Non-linear trend random_oscillations = 0.2 * np.random.normal(size=self.total_length) # Small random noise
# Final sequence: base + non-stationary + trend + noise sequence = base_signal + non_stationary_signal + trend + random_oscillations
# Normalize the sequence sequence = (sequence - np.min(sequence)) / (np.max(sequence) - np.min(sequence)) return sequence
def __len__(self): return self.total_length - self.seq_length_input - self.seq_length_output + 1
def __getitem__(self, idx): input_seq = self.sequence[idx:idx + self.seq_length_input] output_seq = self.sequence[idx + self.seq_length_input:idx + self.seq_length_input + self.seq_length_output] return ( torch.tensor(input_seq, dtype=torch.float32).unsqueeze(-1), torch.tensor(output_seq, dtype=torch.float32).unsqueeze(-1), )
Appendix E Proof

### v2

@@ 1 Introduction
Figure 1: Performance of Mean Squared Error (MSE) on a simple synthetic non-stationary signal.
@@ 2.1 Efficient linear forecasters
Since (Zeng et al., 2023) shows that a simple one-layer linear model can outperforms Transformer forecasters (Zhou et al., 2021; Wu et al., 2021; Zhou et al., 2022) in almost all cases, there has been a rapid emergence of linear forecasters (Oreshkin et al., 2020; Das et al., 2023; Liu et al., 2023) in LTSF. The impressive performance and efficiency continuously challenge this direction. Lin et al. (2024) employs a Linear backbone combined with learnable recurrent cycles to explicitly model periodic structures in time series. Recently, FITS (Xu et al., 2024) introduced a frequency-domain interpolation strategy that utilizes low-pass filters and FFT (Brigham and Morrow, 1967) for time series modeling. With $10k$ parameter, FITS surpassed several existing Transformer models in both inference speed and forecasting performance, establishing itself as a benchmark in the field.
However, FFT assumes that signals are stationary, limiting its ability to capture the temporal localization of transient or non-stationary signals (Liu et al., 2022). Meanwhile, the limited representational capacity of a single-layer linear model typically necessitates a longer look-back window to prevent underfitting and distribution shifts. The parameter count of FITS is primarily determined by the length of the look-back window due to its interpolation-based prediction approach. Consequently, The efficiency of FITS decreases significantly as the lookback window increases.
@@
Figure 2: Overall structure of SWIFT. (i)The DWT module decomposes the time series into two sub-series: the approximation coefficient and the detail coefficient, based on the Haar wavelet; (ii) The convolutional layer is applied for filtering and feature aggregation. (iii)Linear or MLP is used for the mapping of sub-series to make prediction. T denote the length of the look-back window, N is the number of variables (i.e., channels), and S refers to the length of the prediction horizon.
@@ 2.2 DWT method
As a powerful method for time-frequency analysis, DWT is widely used in tasks dealing with time series. (Yang et al., 2022) decomposed the time-series using wavelet decomposition and then utilized CNN and LSTM for prediction. (Zhou et al., 2022) combines Wavelet Transform with frequency enhanced strategy and attention mechanism to capture long range dependencies. (Sasal et al., 2022) utilize a maximal overlap discrete wavelet transformation and build a local transformer model for time-series forecasting. (adawavenet) utilizes learnable lifting-based wavelet transforms to adaptively model non-stationary time series. (wpmixer) leverages level-specific wavelet coefficient decomposition combined with patch-based mixing mechanisms to preserve multi-resolution information.
DWT is also often used for anomaly detection. For instance, (Bhattacharya et al., 2022) used wavelet transform for signal denoising and damage localization. Recently, (Arabi et al., 2024) proposed a method for data augmentation in time-series prediction tasks, which is used to obtain more diverse sequences by eliminating or swapping wavelet coefficients.
@@ 3.1 LTSF problem definition
In multivariate LTSF, time series data contain multiple variables at each time step. Given historical values $\mathcal{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{L_{x}}~{}|~{}\mathbf{x}_{i}\in\mathbb{R}^{d}\}$ where $d$ represents the number of variables and $L_{x}$ represents the length of the look-back window, the objective of LTSF is to predict future values $\mathcal{Y}=\{\mathbf{y}_{1},\ldots,\mathbf{y}_{L_{y}}~{}|~{}\mathbf{y}_{i}\in\mathbb{R}^{d}\}$. The output length $L_{y}$ is usually much longer than the length of the lookback window $L_{x}$, and the feature dimension is not limited to the univariate case ($d\geq 1$).
@@ 3.2 DWT and time-frequency domain
We select Haar because it can make the transform fast and stable, which increases the speed of reasoning across our framework. Although higher-order wavelets such as Daubechies and Symlet generally exhibit better smoothness and frequency localization properties, the Haar wavelet excels at capturing sharp transitions and local discontinuities, which are often critical in time series forecasting tasks involving abrupt changes or short-term pattern shifts. Furthermore, in our implementation of SWIFT, the use of a $1_{st}$-order Haar decomposition (length=2) offers the added advantage of mitigating edge effects in discrete wavelet transforms (DWT) by requiring minimal boundary data. These considerations collectively make the Haar basis particularly well-suited for our application.
@@ 4.1 Structure Overview
We propose the SWIFT which is shown in Figure 2. Firstly, time-series is decomposed by $1_{st}$ order DWT. Then the high-frequency component and the low-frequency component are concatenated and mapped using the same backbone layer after passing through a learnable filter. Finally, the prediction is obtained by performing IDWT on the new components obtained from the mapping.
@@ 4.2 SWIFT Components
After obtaining the representation $\mathbf{Y}$, SWIFT extracts information from this representation by means of convolution and Mapping.
@@ 5.1 Forecasting results
We carefully choose well-acknowledged forecasting models as our benchmark, including (1) Transformer-based methods: FEDformer (Zhou et al., 2022), PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2024). (2) Efficient Linear-based methods: DLinear (Zeng et al., 2023), FITS (Xu et al., 2024), CycleNet (Lin et al., 2024). (3) TCN-based methods: TimesNet (Wu et al., 2022).
@@
Our method is trained with the ADAM optimizer (Kingma and Ba, 2015), using OneCycleLR strategy (Paszke et al., 2019) to adjust the learning rate. For the kernel size of the filter, we choose suitable size in {3, 9, 13, 17}. We evaluate all models across prediction horizons of {96, 192, 336, 720}. For historical input lengths, we follow (Xu et al., 2024) by treating $T$ as a hyperparameter, systematically conducting grid search within {96, 180, 360, 720} to identify the optimal length for each model. This approach accounts for observations that some models degrade with longer histories (e.g., iTransformer (Liu et al., 2024) on ETT datasets (Xu et al., 2024)). Our model (SWIFT) and CycleNet (Lin et al., 2024) use a lookback window length of 720 because both of them achieve best performance at $T$=720, as their accuracy improves with extended input lengths. For all other models, we fix other hyperparameters to their original settings from official implementations, except for historical input lengths. We rerun all the experiment with code and script provided by their official implementation.
@@
Comprehensive forecasting results are listed in Table 1 and Table 2 with the best in bold and the second underlined. The lower MSE indicates the more accurate prediction result. As shown in table 1 and table 2, SWIFT performs well in the forecasting task. Overall, SWIFT achieves state of the art performance. Due to the nonlinear mapping capability of MLP compared to Linear, the MLP version of SWIFT performs better on high-dimensional datasets such as Electricity and Solar-Energy (i.e., datasets with more than 100 channels). In summary, SWIFT achieves comparable or even superior performance in all 7 datasets with even a very simple Linear or MLP, while requiring nearly $100K\times$ fewer parameters than existing methods (i.e., Transformer-based methods).
Table 3: Number of trainable parameters, MACs, and training time of models with less than 1M parameters, under look-back window=720 and forecasting horizon=96 on Electricity.
@@
Table 4: Ablation study 1 of SWIFT. For dataset etth1 and ettm1, we choose Linear version of SWIFT; for dataset traffic, we use MLP version of SWIFT.
@@
Table 5: Ablation study 2 of SWIFT. IMP is the improvement between with $Share$ and $Split$ result, where a larger value indicates a better improvement. $Share$ corresponds to the use of only one linear layer, while $Split$ corresponds to the use of one linear layer for each frequency band. The model we used is SWIFT / Linear.
@@
Table 3 presents the number of trainable parameters and MACs for various linear-based time series forecasting (TSF) models using a look-back window of 720 and a forecasting horizon of 96 on the Electricity dataset. The table clearly demonstrates the exceptional efficiency of SWIFT compared to other models. Among all efficient models, SWIFT stands out with significantly fewer parameters and much faster training times. SWIFT-Linear requires only 15% of the FITS parameters and 60% of its MACs, while achieving comparable or even superior performance to these state-of-the-art efficient forecasting models. It should be noted that SWIFT-Linear’s parameter count is also much lower than Dlinear, which has 138.4K parameters. Moreover, while the parameter count of FITS increases rapidly when using longer look-back windows for forecasting, SWIFT does not exhibit this issue.
@@ 6 Conclusion and Future work
Appendix A Proof
@@
Appendix B More results
B.1 Ablations on Channel Independence
To investigate the impact of Channel Independence (CI) on model performance, we conduct an ablation study comparing SWIFT models with and without the CI mechanism across four ETT benchmark datasets and Traffic dataset. Table 7 presents the Mean Squared Error (MSE) results for different prediction horizons.
Table 7: Ablation results on Channel Independence in SWIFT Model Performance (MSE) across ETT and Traffic.
Dataset Model L=96 L=192 L=336 L=720
ETTh1 Not $CI$ 0.367 0.395 0.420 0.430
$CI$ 0.375 0.413 0.430 0.450
ETTh2 Not $CI$ 0.268 0.329 0.351 0.383
$CI$ 0.293 0.337 0.365 0.398
ETTm1 Not $CI$ 0.307 0.336 0.364 0.413
$CI$ 0.300 0.335 0.368 0.420
ETTm2 Not $CI$ 0.161 0.214 0.267 0.348
$CI$ 0.161 0.215 0.268 0.348
Traffic Not $CI$ 0.368 0.382 0.396 0.430
$CI$ 0.425 0.444 0.462 0.487
  • Note: All experiments use sequence length = 720 and seed = 2023.
  • Bold: Best performance. underline: Performance of the tie.
B.2 Ablations on Wavelets
To investigate the impact of different wavelet filters on SWIFT model performance, we conduct a comprehensive ablation study comparing three representative wavelets: Haar, Daubechies-2 (DB2), and Symlet-4 (Sym4) across multiple benchmark datasets. The choice of wavelet filter is crucial for the SWIFT architecture, as it directly affects both computational efficiency and forecasting accuracy.
Table 8 presents the comparative results of MSE performance across different prediction horizons. The experimental findings reveal that Haar wavelets consistently achieve the best performance across all datasets and prediction lengths, demonstrating universally superior forecasting accuracy. Specifically, Haar wavelets outperform both DB2 and Sym4 wavelets with MSE improvements ranging from 2.7% to 5.7% on ETTh1 compared to the second-best performer. On ETTm1, Haar maintains consistent advantages with improvements of 0.6% to 2.8% over alternative wavelets.
Most notably, the Weather dataset shows the most significant performance gains with Haar wavelets, achieving improvements of 7.3-8.5% (L=96), 5.7% (L=192), 3.7-4.1% (L=336), and 1.9-2.5% (L=720) compared to DB2 and Sym4. Interestingly, DB2 and Sym4 wavelets exhibit similar performance patterns, with Sym4 showing marginal improvements over DB2 in some configurations, but neither approaching the consistent superiority of Haar wavelets across all experimental scenarios.
Beyond forecasting accuracy, the computational efficiency considerations strongly favor Haar wavelets. The Haar wavelet filter possesses the shortest possible support length among orthogonal wavelets, consisting of only two coefficients [1, -1]. This minimal filter length is particularly crucial for our proposed sharing strategy, which constitutes the core mechanism enabling SWIFT’s computational efficiency. Longer wavelet filters, such as DB2 with its four-coefficient structure and Sym4 with its eight-coefficient structure, introduce additional computational overhead and memory requirements that are incompatible with our efficient sharing framework. The combination of superior forecasting performance and optimal computational characteristics makes Haar wavelets the ideal choice for the SWIFT architecture. The results demonstrate that Haar filters not only maintain competitive accuracy but also provide the necessary efficiency gains that enable real-time processing capabilities essential for practical time series forecasting applications.
Table 8: Ablation Study on Different Wavelets in SWIFT Model Performance (MSE).
Dataset Wavelet L=96 L=192 L=336 L=720
ETTh1 Haar 0.367 0.395 0.420 0.430
DB2 0.377 0.413 0.444 0.456
Sym4 0.374 0.412 0.447 0.462
Coif 0.375 0.430 0.450 0.459
ETTm1 Haar 0.307 0.336 0.364 0.413
DB2 0.309 0.343 0.373 0.418
Sym4 0.310 0.339 0.365 0.415
Coif 0.320 0.341 0.368 0.415
Weather Haar 0.140 0.183 0.235 0.307
DB2 0.153 0.194 0.245 0.315
Sym4 0.151 0.194 0.244 0.313
Coif 0.152 0.195 0.243 0.315
  • Note: All experiments use sequence length = 720 and seed = 2023.
  • Bold: Best performance for each prediction length.